# Solving the Bongard-logo Problem by Modeling a Probabilistic Model

### v7

@@ III-B PMoC(Probability Model of Concept)
After conducting experiments, we discovered that SBSD had achieved limited progress in inference accuracy. However, this paper aims to make more contributions and breakthroughs. Therefore, by carrying forward the idea of analyzing primary and auxiliary group distributions in SBSD, this paper designs the deep model PMoC (Probability Model of Concept). PMoC explore an alternative approach by shifting the constraint conditions. The training objective has been shifted from focusing on the deep model’s ability to disentangle distribution $p_{i}^{\prime}(z|y)$ and distribution $q_{i}^{\prime}(z|y)$ to emphasizing its capacity to compute the probability of a representation under the distribution $p_{i}^{\prime}(z|y)$. For this purpose, we assume that $p_{i}^{\prime}(z|y)$ follows a multivariate Gaussian distribution.
Our PMoC consists of two components: module $f_{\theta}(z|x)$ and module $g_{\omega}(\mu,\sigma^{2}|z)$. The module $f_{\theta}(z|x)$ maps the input $x$ to its latent representation $z$, while $g_{\omega}(\mu,\sigma^{2}|z)$ is utilized to compute the $\mu$ and $\sigma^{2}$ for the Gaussian distribution $p^{\prime}_{i}(z|y)$.
In this paper, we define $\{x_{ij}\}_{j=1}^{6}$ as the primary group samples for the $i$-th Bongard-Logo question, with $j$ denoting the sample number. Similarly, $\{x_{ij}\}_{j=8}^{13}$ represents the auxiliary group samples for the $i$-th question. $x_{i7}$ and $x_{i14}$ are the representations of the samples to be classified. Meanwhile, $\{z_{ij}\}_{j=1}^{14}$ represents the encoded Bongard-Logo samples $\{x_{ij}\}_{j=1}^{14}$, and the encoding process is implemented by $f_{\theta}(z_{ij}|x_{ij})$. Furthermore, we utilize a deep network, denoted as $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$, to fit an optimizable multivariate Gaussian distribution $p^{\prime}_{i}(z|\{z_{ij}\}^{6}_{j=1},\omega)$, based on the primary group samples features $\{z_{ij}\}^{6}_{j=1}$. Subsequently, we calculate the probability of each representation in $\{z_{ij^{\prime}}\}_{j^{\prime}=7}^{14}$ belonging to this distribution, denoted as $\{p^{\prime}_{i}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$. The network $f_{\theta}(z_{ij}|x_{ij})$ maps the input $x_{ij}$ to its latent representation $z_{ij}$, while $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is utilized to compute the distribution of the latent representations $\{z_{ij}\}^{6}_{j=1}$, and the distribution is modeled as Gaussian distribution. In other words, $f_{\theta}(\cdot|\cdot)$ serves as an image encoder, and $g_{\omega}(\cdot|\cdot)$ as a logical distribution fitter. Thus, we employ a Convolutional Neural Network (CNN) to fit $f_{\theta}(\cdot|\cdot)$ and a Transformer-Encoder to fit $g_{\omega}(\cdot|\cdot)$. Further design details are provided as follows.
@@
The network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is much more of its name. It encodes the mean and variance of the $i$-th primary group distribution based on the latent representations of the $i$-th primary group samples, and $i$ still represents the problem index in Bongard-Logo. Specifically, we uniformly process each perspective $z_{ij}$ within $z^{\prime}_{ij}$. Under a single perspective, we treat $\{z_{ij}\}^{6}_{j=1}$ as six tokens and, after appending two optimizable vectors along with positional embedding, input them into $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ with a Transformer-Encoder backbone for computing attention results. We extract two optimizable vectors from the attention results, serving as the mean $\mu_{i}\in R^{d}$ and logarithm of variance $\log(\sigma_{i}^{2})\in R^{d}$ for a multivariate Gaussian distribution. The encoding distribution pattern of module $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ follows a similar approach to that of VAE[7]; however, it does not incorporate variational inference in its design. The aforementioned process parameterizes $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and outlines the computation of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$. We feed 8 vectors $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ into this distribution for calculating 8 logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$ in one perspective. In this paper, we calculate the final probability ${p}_{i}^{\prime}(z^{\prime}_{ij^{\prime}})$ by averaging the logical probabilities ${p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)$ obtained from all perspectives. By utilizing the cross entropy loss function, we constrain the final probability ${p}_{i}^{\prime}(z^{\prime}_{i7})$ to approach 1 while ensuring other probabilities $\{{p}_{i}^{\prime}(z^{\prime}_{ij})\}_{j=8}^{14}$ are close to 0. Figure 6 illustrates in detail the forward process of the network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ under a single perspective in $z^{\prime}_{ij}\in R^{n\times d}$. However, this approach did not yield satisfactory results.
Figure 6: Feedforward process of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$.
@@
To counter such ineffectiveness, in this paper, we train a deep network to directly fit probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Inevitably the network’s formulation has been transformed from $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ to $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$, thereby simplifying the problem. Different from previous approaches $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ that fit the distribution $p_{i}^{\prime}(z|y)$, this paper employs a Transformer-Encoder to fit the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$ directly, which enables the direct computation of the probability that $z_{ij^{\prime}}$ belongs to the distribution $p_{i}^{\prime}(z|y)$.
@@
For the implementation on the new form of $g_{\omega}(\cdot|\cdot)$, the function $g_{\omega}(\cdot|\cdot)$ maintains the egalitarian approach in processing each perspective $z_{ij}\in R^{d}$ within the outcomes $z^{\prime}_{ij}\in R^{n\times d}$ encoded by $f_{\theta}(z_{ij}|x_{ij})$. Within the confines of a singular perspective, we amalgamate $\{z_{ij}\}^{6}_{j=1}$ with one of the pending representations $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, resulting in a set of seven tokens. Following the embedding of positional encoding, these tokens are then systematically fed into a Transformer-Encoder, aimed at yielding 7 attention results. All 7 attention results are mapped to 7 probabilities by a MLP. The mean of the mapped outcomes is considered as the logical probability ${p}_{i}^{\prime}(z_{ij^{\prime}}|y)$ corresponding to the specified perspective. This constitutes the feedforward process of the $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$. Each pending representation $z_{ij^{\prime}}$ in $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, receives a corresponding probability output $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Each logical probability stored in $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$ is obtained based on a single perspective. By averaging the logical probabilities across all perspectives, we derive the final probability values $\{p_{i}^{\prime}(z^{\prime}_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$. In this paper, we still optimize the model using the cross-entropy loss function. Figure 7 visually depicts the process of $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$ handling a single perspective.
@@ IV Methodology for abstract reasoning
Previous research has demonstrated limited success in learning abstract concepts related to positional relationships among entities, particularly when addressing the Bongard-Logo and RPM problems. This limitation is evident in challenges posed by the OIG and D-9 problem sets from RAVEN, as well as the ff and hd issues of Bongard-Logo and the Neurtal problem in PGM. The RS-model[31], however, overcame these difficulties by explicitly incorporating additional structural cues (meta data, auxiliary label), achieving commendable inferential accuracy.
In this section, we introduce the Pose-Transformer, an enhancement to the Transformer-Encoder framework that incorporates the pose matrix from the Capsule Network[33]. This integration implicitly enhances the network’s ability to discern positional relationships among local representations, thereby bolstering its learning capabilities on graphical abstract reasoning problem. Furthermore, we have lightweighted Pose-Transformer into Straw-Pose Transformer, which achieves a lightweight version of the former while maintaining its effectiveness. We present and delve into the intricacies of our proposed Pose-Transformer, discussing its potential to revolutionize learning in tasks that involve complex positional relationships through experiments.
@@ IV-A Pose-tansformer
We delve into the intricacies of the Transformer-Encoder framework, particularly when it processes logical inputs to generate attention-weighted outputs. During this process, all output tokens become associated with varying degrees of global information derived from the logical inputs[27, 34]. Operating on a multi-head self-attention mechanism, the Transformer-Encoder segments the input tokens into multiple sub-vectors and computes their self-attention results in parallel[34]. From an informational standpoint, we postulate that these segmented sub-vectors possess lesser information compared to the holistic tokens encoding global information, effectively rendering them as representations of localized information. Consequently, the relative positional relationships among these localized informational units warrant a summarization and analysis.
Specifically, we have designed a novel fundamental module termed “pose matrix embedding” tailored for the RPM and Bongard-Logo problems. When logical inputs $\{\tilde{V}_{i}|i\in[1,n],\tilde{V}_{i}\in R^{d}\}$ pass through a single Transformer-Encoder layer, this paper segments the $n$ output tokens $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$, attention results, into $m$ local vectors $\{H_{ij}|i\in[1,n],j\in[1,m],H_{ij}\in R^{d/m}\}$ of the attention head size. A set of mapping matrices $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$ is established for each local vector, with the number of matrices $k$ within a set being identical to the token count $n$ of the logical input. These mapping matrices facilitate the transformation of local vectors into pose vectors $\left(H_{ij}\times W_{ijk}\right)$. We compute the entire set of pose vectors $\{P_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],P_{ijk}\in R^{d}\}$ resulting from the mapping of each local vector. By summing and squashing these pose vectors and adding them to $\{V_{i}|i\in[1,n],Vi\in R^{d}\}$, we obtain the logical output $\{O_{k}|k\in[1,n],O_{k}\in R^{d}\}$ of the foundational block of Pose-Transformer. The squashing function is formulated as shown in the equation below:
@@
In the formula, the term “segment” represents the aforementioned cutting process. The process of feeding the input vector set into the transformer-encoder to obtain the output is denoted as “T-Encoder”.
@@
This fundamental block is followed by a feedforward block, forming the basic building block of the Pose-Transformer. The design of the feedforward block is consistent with that in the Transformer-Encoder architecture. Notably, the proposed fundamental block is dimension-preserving, allowing for the stacking of multiple blocks to adapt to different scales of graphical reasoning problems. The Pose-Transformer with $N$ logical layers is depicted in the figure 9. When $N=1$, the Pose-Transformer collapses into a regular Transformer-Encoder . In Capsule Networks, similarity weights are distributed across higher-level capsules and computed in a bottom-up fashion, whereas attention weights in the Transformer-Encoder are distributed over lower-level representations and computed top-down. In the Pose-Transformer with $N>1$, similarity weights for local representations can be calculated at multiple levels and scales from the outside in. This is ensured by the fact that each logical layer in the Pose-Transformer is succeeded by a standard Transformer-Encoder layer, which can be present in the subsequent logical layer or the terminating layer, responsible for recalculating the similarity weights.

### v8

@@ III-B PMoC(Probability Model of Concept)
After conducting experiments, we discovered that SBSD had achieved limited progress in reasoning accuracy. However, this paper aims to make more contributions and breakthroughs. Given the difficulties associated with utilizing the SBSD approach, which involves optimizing deep networks based on the distances between distributions, this paper proposes the design of the PMoC (Probability Model of Concept) network. PMoC explores an alternative approach by shifting the constraint conditions. The training objective shifts from focusing on the deep model’s ability to disentangle distribution $p_{i}^{\prime}(z|y)$ and distribution $q_{i}^{\prime}(z|y)$ to emphasizing its capacity to compute the probability of a sample belonging to the distribution $p_{i}^{\prime}(z|y)$. For this purpose, we assume that $p_{i}^{\prime}(z|y)$ follows a multivariate Gaussian distribution.
From a macroscopic perspective, our PMoC framework consists of two components: module $f(z|x)$ and module $g(\mu,\sigma^{2}|z)$. We leverage the $f(z|x)$ module to map Bongard-Logo images into latent representations, and the $g(\mu,\sigma^{2}|z)$ module to deduce a Gaussian distribution based on the representations of samples from the primary group. Subsequently, we compute the probabilities of representations from the auxiliary group and the test samples belonging to this distribution. Finally, we employ a cross-entropy loss function to constrain these probabilities, optimizing both the mapping capability of the $f(z|x)$ module and the distribution computation ability of the $g(\mu,\sigma^{2}|z)$ module.
Specifically, we utilize a deep network $f_{\theta}(z_{ij}|x_{ij})$ (serving as the $f(z|x)$ module) to transform the Bongard-Logo images $\{x_{ij}\}^{14}_{j=1}$ into their corresponding latent representations $\{z_{ij}\}^{14}_{j=1}$. Then, another deep network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ (acting as the $g(\mu,\sigma^{2}|z)$ module) is employed to derive an optimizable multivariate Gaussian distribution $p^{\prime}_{i}(z|\{z_{ij}\}^{6}_{j=1},\omega)$ based on the latent representations of the primary group samples $\{z_{ij}\}^{6}_{j=1}$. Next, we compute the probabilities of each latent representation $z_{ij^{\prime}}$ (where $j^{\prime}$ ranges from 7 to 14) belonging to this distribution, denoted as the set $\{p^{\prime}_{i}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$. In essence, the network $f_{\theta}(z_{ij}|x_{ij})$ functions as an image encoder, mapping input images to their latent representations, while $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ serves as a logical distribution fitter, calculating the distribution of latent representations and modeling it as a Gaussian distribution. To this end, we adopt a Convolutional Neural Network (CNN) to fit $f_{\theta}(z_{ij}|x_{ij})$ and a Transformer-Encoder to fit $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$. Further details regarding the design are provided below. This framework is visualized in Figure 5.
Figure 5: Framework of PMoC.
@@
The network $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ is much more of its name. It encodes the mean and variance of the $i$-th primary group distribution based on the latent representations of the $i$-th primary group samples, and $i$ still represents the problem index in Bongard-Logo. Specifically, we uniformly process each perspective $z_{ij}$ within $z^{\prime}_{ij}$. Under a single perspective, we treat $\{z_{ij}\}^{6}_{j=1}$ as six tokens and, after appending two optimizable vectors along with positional embedding, input them into $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ with a Transformer-Encoder backbone for computing attention results. We extract two optimizable vectors from the attention results, serving as the mean $\mu_{i}\in R^{d}$ and logarithm of variance $\log(\sigma_{i}^{2})\in R^{d}$ for a multivariate Gaussian distribution. The encoding distribution pattern of module $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ follows a similar approach to that of VAE[7]; however, it does not incorporate variational inference in its design. The aforementioned process parameterizes $p_{i}^{\prime}(z|y)$ as a multivariate Gaussian distribution and outlines the computation of $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ and this process can be expressed as Figure 7.
Figure 7: Parameterizes process of $p_{i}^{\prime}(z|y)$
We feed 8 vectors $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ into this distribution for calculating 8 logical probabilities $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)\}_{j^{\prime}=7}^{14}$ in one perspective. In this paper, we calculate the final probability ${p}_{i}^{\prime}(z^{\prime}_{ij^{\prime}})$ by averaging the logical probabilities ${p}_{i}^{\prime}(z_{ij^{\prime}}|\{z_{ij}\}^{6}_{j=1},\omega)$ obtained from all perspectives. The calculation process of the final probability can be expressed as Figure 8.
Figure 8: The calculation process of the final probability
By utilizing the cross entropy loss function, we constrain the final probability ${p}_{i}^{\prime}(z^{\prime}_{i7})$ to approach 1 while ensuring other probabilities $\{{p}_{i}^{\prime}(z^{\prime}_{ij})\}_{j=8}^{14}$ are close to 0. Figure 9 illustrates in detail the forward process of the PMoC. However, this approach did not yield satisfactory results.
Figure 9: Feedforward process of PMoC.
@@
To counter such ineffectiveness, we decide to train a deep network to directly fit probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. We argue that, rather than laboriously analyzing and fitting a distribution, it is more efficient to directly compute the probability of a sample belonging to that distribution. Inevitably the network’s formulation has been transformed from $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ to $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$, thereby shifting the problem. Different from previous approaches $g_{\omega}(\mu_{i},\sigma_{i}|\{z_{ij}\}^{6}_{j=1})$ that fit the distribution $p_{i}^{\prime}(z|y)$, this paper employs a Transformer-Encoder to fit the probability $p_{i}^{\prime}(z_{ij^{\prime}}|y)$ directly, which enables the direct computation of the probability that $z_{ij^{\prime}}$ belongs to the distribution $p_{i}^{\prime}(z|y)$.
For the implementation on the new form of $g_{\omega}(\cdot|\cdot)$, the function $g_{\omega}(\cdot|\cdot)$ maintains the egalitarian approach in processing each perspective $z_{ij}\in R^{d}$ within the outcomes $z^{\prime}_{ij}\in R^{n\times d}$ encoded by $f_{\theta}(z_{ij}|x_{ij})$. We amalgamate $\{z_{ij}\}^{6}_{j=1}$ with one of the pending representations $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$, resulting in a set of seven tokens. Following the embedding of positional encoding, these tokens are then systematically fed into a Transformer-Encoder, aimed at yielding 7 attention results. All 7 attention results are mapped to 7 probabilities by a MLP. The mean of the mapped outcomes is considered as the logical probability ${p}_{i}^{\prime}(z_{ij^{\prime}}|y)$ corresponding to the specified perspective. This constitutes the feedforward process of the $g_{\omega}(p_{i}^{\prime}(z_{ij^{\prime}}|y)|\{z_{ij}\}^{6}_{j=1},z_{ij^{\prime}})$. This feedforward process can be expressed as Figure 10.
@@
Each pending representation $z_{ij^{\prime}}$ within $\{z_{ij^{\prime}}\}^{14}_{j^{\prime}=7}$ receives a corresponding probability output $p_{i}^{\prime}(z_{ij^{\prime}}|y)$. Each logical probability stored in $\{{p}_{i}^{\prime}(z_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$ is obtained based on a single perspective. By averaging the logical probabilities across all perspectives, we derive the final probability values $\{p_{i}^{\prime}(z^{\prime}_{ij^{\prime}}|y)\}_{j^{\prime}=7}^{14}$. The new calculation process of the final probability can be expressed as Figure 11.
Figure 11: The new calculation process of the final probability
In this paper, we still optimize the model using the cross-entropy loss function. It is worth mentioning that the essence of the final version of PMoC lies in its function as a distribution distance metric, thereby necessitating the satisfaction of Lipschitz continuity. To ensure this, we have imposed spectral normalization constraints on the parameters of PMoC. Figure 12 visually depicts the process of PMoC.
Figure 12: Feedforward process of PMoC 2.0 ver.
@@
Despite the apparent disadvantage of probabilistic clustering reasoning solvers, we persist in designing and developing a Bongard-Logo solver in the form of a probabilistic model. Our aim is not to rival the most advanced solutions available, but rather to contribute to the diversity of deep learning research outcomes. We are motivated by the goal of fostering innovation and exploration beyond the mainstream, encouraging a wider range of approaches and perspectives in the field. We are confident that a probabilistic clustering approach can offer more intuitive probability assessments for clustering decisions, which could be critically important in certain analogous fields.
@@ IV Methodology for abstract reasoning
Previous research has demonstrated limited success in learning abstract concepts related to positional relationships among entities, particularly when addressing the Bongard-Logo and RPM problems. This limitation is evident in challenges posed by the OIG and D-9 problem sets from RAVEN, as well as the ff and hd issues of Bongard-Logo and the Neurtal problem in PGM.
In this section, we introduce the Pose-Transformer, an enhancement to the Transformer-Encoder framework that incorporates the pose matrix from the Capsule Network[32]. This integration implicitly enhances the Transformer-Encoder’s ability to discern positional relationships among local representations, thereby bolstering its learning capabilities on visual abstract reasoning problem. Furthermore, we have lightweighted Pose-Transformer into Straw-Pose-Transformer, which achieves a lightweight version of the former while maintaining its effectiveness. We present and delve into the intricacies of our proposed Pose-Transformer, discussing its potential to revolutionize learning in tasks that involve complex positional relationships through experiments.
@@ IV-A Pose-tansformer
We delve into the intricacies of the Transformer-Encoder framework, particularly when it processes logical inputs to generate attention-weighted outputs. During this process, all output tokens become associated with varying degrees of global information derived from the logical inputs[26, 33]. Operating on a multi-head self-attention mechanism, the Transformer-Encoder segments the input tokens into multiple sub-vectors and computes their self-attention results in parallel[33]. From an informational standpoint, we posit that these segmented sub-vectors possess lesser information compared to the holistic tokens encoding global information, effectively rendering them as representations of localized information. Consequently, the relative positional relationships among these localized informational units warrant a summarization and analysis.
Notably, in Capsule Networks[32], a set of pose mapping matrix is set to map local information features into pose vectors, and the Gaussian routing algorithm is employed to calculate the pose consistency between these pose vectors, thereby gauging the relative positional relationship among local features[32]. This paper proposes that a similar pose mapping process could be integrated into the Transformer-Encoder, enabling it to be more sensitive to the relative positional relationships among local features when handling visual tasks. Therefore, we have designed Pose-Transformer and have high expectations for its performance on visual abstract reasoning problems.
Specifically, we have designed a novel fundamental module termed “pose matrix embedding”. When logical inputs $\{\tilde{V}_{i}|i\in[1,n],\tilde{V}_{i}\in R^{d}\}$ pass through a single Transformer-Encoder layer, this paper segments the $n$ output tokens $\{V_{i}|i\in[1,n],V_{i}\in R^{d}\}$, attention results, into $m$ local vectors $\{H_{ij}|i\in[1,n],j\in[1,m],H_{ij}\in R^{d/m}\}$ of the attention head size. A set of mapping matrices $\{W_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],W_{ijk}\in R^{(d/m)\times d}\}$ is established for each local vector, with the number of matrices $k$ within a set being identical to the token count $n$ of the logical input. These mapping matrices facilitate the transformation of local vectors into pose vectors $\left(H_{ij}\times W_{ijk}\right)$. We compute the entire set of pose vectors $\{P_{ijk}|i\in[1,n],j\in[1,m],k\in[1,n],P_{ijk}\in R^{d}\}$ resulting from the mapping of each local vector. By summing and squashing these pose vectors and embedding them to $\{V_{i}|i\in[1,n],Vi\in R^{d}\}$, we obtain the logical output $\{O_{k}|k\in[1,n],O_{k}\in R^{d}\}$ of the foundational block of Pose-Transformer. The squashing function is formulated as shown in the equation below:
@@
In the formula, the operator “segment” represents the aforementioned segment process. The process of feeding the input vector set into the Transformer-Encoder to obtain the attention results is denoted as “T-Encoder”.
@@
This fundamental block is followed by a feedforward block, forming the basic building block of the Pose-Transformer. The design of the feedforward block is consistent with that in the Transformer-Encoder architecture. Notably, the proposed fundamental block is dimension-preserving, allowing for the stacking of multiple blocks to adapt to different scales of visual reasoning problems. The Pose-Transformer with $N$ logical layers is depicted in the figure 14. When $N=1$, the Pose-Transformer collapses into a regular Transformer-Encoder . In Capsule Networks, similarity weights are distributed across higher-level capsules and computed in a bottom-up fashion, whereas attention weights in the Transformer-Encoder are distributed over lower-level representations and computed top-down. In the Pose-Transformer with $N>1$, similarity weights for local representations can be calculated at multiple levels and scales from the outside in. This is ensured by the fact that each logical layer in the Pose-Transformer is succeeded by a standard Transformer-Encoder layer, which can be present in the subsequent logical layer or the terminating layer, responsible for recalculating the similarity weights.